OpenStudy (anonymous):

find the length of JK with endpoints J(-3,25) and K(-7,28) express your answer in simplest radical form

OpenStudy (anonymous):

sqrt ((28-25)^2 + (-7--3)^2)) = sqrt (9+ 16) = sqrt (25) = 5